OpenStudy (anonymous):

Can you guys help me out with this one problem medal will be given factor m6-n6

OpenStudy (unklerhaukus):

\[m^6-n^6\] \[m^3=M\\n^3=N\] \[M^2-N^2=(M+N)(M-N)\]

OpenStudy (perl):

\[(m-n)(m+n)(m^2+mn+n^2)(m^2-mn+n^2)\]

OpenStudy (anonymous):

which one is right?

OpenStudy (perl):

me

OpenStudy (unklerhaukus):

both are equivalent, however i think my way leads to a more factored form

OpenStudy (anonymous):

ok then thanks to both of you guys.

OpenStudy (perl):